OpenStudy (anonymous):

∫(1/sqrt(1-4x-x^2))dx

OpenStudy (anonymous):

I got as far as completing the square

OpenStudy (anonymous):

This requires a trig substitution.

OpenStudy (anonymous):

Or using a table.

OpenStudy (anonymous):

continueee...

OpenStudy (anonymous):

What do you have when you completed the square?

OpenStudy (anonymous):

1/Sqrt(1-(-4+4+4x+x^2)) = 1+4-(4++x+x^2) => 5-(x+2)^2

OpenStudy (anonymous):

Good.

OpenStudy (anonymous):

Now we need to let \(\sqrt{5}sin(\theta ) = x + 2\)

OpenStudy (zarkon):

complete the square on \[1-4x-x^2\] then use the substitution \[\sqrt{5}u=x+2\]

OpenStudy (anonymous):

Even easier, do a u-sub first. Let u=x+2 du=dx Then you have: \[\int\limits \frac{du}{\sqrt{5-u^2}}\]

OpenStudy (anonymous):

Then use the substitution: \[\sqrt5\sin(\phi)=u\] \[du=\sqrt5 \cos(\phi)d \phi\] \[\sqrt5\cos(\phi)=\sqrt{5-u^2}\] Then you have: \[\int\limits \frac{\sqrt5 \cos(\phi) d \phi}{\sqrt5 \cos(\phi)}=\int\limits d \phi=\phi+C\] So solve for phi. \[\sqrt5 \sin(\phi)=u \rightarrow \phi=\arcsin(\frac{u}{\sqrt5})\] But u=x+2 \[\phi=\arcsin(\frac{x+2}{\sqrt5})\] So your integral is: \[\arcsin(\frac{x+2}{\sqrt5})+C\]